OpenStudy (ab123):

I need to know the steps of how to do this. The midpoint of line CD is E(-1,0). One endpoint is C(5,2). What are the coordinates of the other endpoint?

OpenStudy (danjs):

The midpoint is half way along the line from C to D. 

OpenStudy (danjs):

You can look at it and see how much each coordinate changes.. from C to E, and this will be the same change from E to D... From C to E, x changes -6 y changes -2 The same changes happen from E to D, take off 6 on the x and 2 on the y D(x,y) = (-1-6 , 0 - 2) = (-7 , -2)

OpenStudy (danjs):

Or if you like the formulas, the midpoint M(x,y) between points (x1,y1) and (x2,y2) is \[\large M(x,y) = (\frac{ x _{1}+x _{2} }{ 2 } , \frac{ y _{1}+y _{2} }{ 2 })\]

OpenStudy (ab123):

How would I do it using the formula?

OpenStudy (danjs):

They give you one endpoint and the midpoint. C(x,y) = (5,2) M(x,y) = (-1,0) \[(-1,0) = (\frac{ 5 + x _{2} }{ 2 }, \frac{ 2 + y _{2} }{ 2 })\]

OpenStudy (danjs):

The other endpoint (x2,y2) can be found from that... \[\large -1 = \frac{ 5+x _{2} }{ 2 } ~~~~~and~~~~~0=\frac{ 2+y _{2} }{ 2 }\]

OpenStudy (danjs):

That gives you the other endpoint D \[x _{2}=-7~~~~~and~~~~~y _{2}=-2\] D(-7,-2)

OpenStudy (ab123):

Can you explain how you got -7 from \[\frac{ 5+x _{2} }{ 2 }\]

OpenStudy (ben00):

Well that's complicated.

OpenStudy (ab123):

@DanJS ?

OpenStudy (danjs):

The midpoint formula is \[\large M(x,y) = (\frac{ x _{1}+x _{2} }{ 2 } , \frac{ y _{1}+y _{2} }{ 2 })\] that means the x-coordinate of midpoint M is half the sum of the x coordinates x1 and x2 of the two endpoints. The same for the y coordinates.

OpenStudy (danjs):

The midpoint is M(-1,0) The given end point is C(5,2) Putting that into the midpoint formula \[\large M(-1,0) = (\frac{ 5+x _{2} }{ 2 } , \frac{ 2+y _{2} }{ 2 })\]

OpenStudy (danjs):

So you get matching the x and y coordinates the two equations. \[-1=\frac{ 5+x _{2} }{ 2 }~~~~~and~~~~~0=\frac{ 2+y _{2} }{ 2 }\]

OpenStudy (ab123):

yes, but how did you get from that to (-7,-2)

OpenStudy (danjs):

That is the point (x2,y2) you are finding... Solve the two equations \[-1=\frac{ 5+x _{2} }{ 2 }\] \[-2=5+x _{2}\] \[x _{2}=-7\]

OpenStudy (ab123):

Ok, thanks for your patience.